# Core equality of real sequences

### v1

@@ 3. Proofs
where $\tilde{\mu}^{\star}(X):=\limsup_{n}\tilde{A}^{+}\bm{1}_{X}$ for each $X\subseteq S$. Lastly, observe that $\tilde{A}^{+}$ has the following properties: $\limsup_{n}\sum_{k}\tilde{a}_{n,k}^{+}=1$, $\lim_{n}\tilde{a}_{n,k}^{+}=0$, and $\sum_{k}\tilde{a}_{n,k}^{+}<\infty$ for all $n,k\in\omega$. In particular, even if $\tilde{A}^{+}$ is not necessarily regular, it satisfies the hypotheses of [8, Theorem 6.2 and Proposition 7.2]. Hence, we proceed verbatim as in Claim 4 and we obtain that $\tilde{\mathcal{I}}$ is not tall. Since $S$ is arbitrary, we conclude that $\mathcal{I}$ is a Fréchet ideal. ∎
Thanks to Claims 1, 3, and 5 we know that $\mathcal{I}$ is an analytic $P$-ideal which is also Fréchet. Then, it is known that $\mathcal{I}$ is necessarily $\mathrm{Fin}$ or $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ or ×$\emptyset\times\mathrm{Fin}$, see e.g. [9, Corollary 1.2.10] or [11, Theorem 2.26]. Finally, $\mathcal{I}$ has to be also a $P^{+}$-ideal by Claim 2. Hence, it is immediate to check that ×$\mathcal{I}\neq\emptyset\times\mathrm{Fin}$. This concludes the proof. ∎

### v2

@@ 3. Proofs
where $\tilde{\mu}^{\star}(X):=\limsup_{n}\tilde{A}^{+}\bm{1}_{X}$ for each $X\subseteq S$. Lastly, observe that $\tilde{A}^{+}$ has the following properties: $\limsup_{n}\sum_{k}\tilde{a}_{n,k}^{+}=1$, $\lim_{n}\tilde{a}_{n,k}^{+}=0$, and $\sum_{k}\tilde{a}_{n,k}^{+}<\infty$ for all $n,k\in\omega$. In particular, even if $\tilde{A}^{+}$ is not necessarily regular, it satisfies the hypotheses of [8, Theorem 6.2 and Proposition 7.2]. Hence, we proceed verbatim as in Claim 4 and we obtain that $\tilde{\mathcal{I}}$ is not tall. Since $S$ is arbitrary, we conclude that $\mathcal{I}$ is a nowhere tall ideal. ∎
Thanks to Claims 1, 3, and 5 we know that $\mathcal{I}$ is an analytic $P$-ideal which is also nowhere tall (notice that such properties are invariant under isomorphisms). Then, it is known that $\mathcal{I}$ is necessarily $\mathrm{Fin}$ or (an isomorphic copy on $\omega$ of) $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ or {}×$\{\emptyset\}\times\mathrm{Fin}$, see e.g. [9, Corollary 1.2.11] or [11, Theorem 2.26]. Finally, $\mathcal{I}$ has to be also a $P^{+}$-ideal by Claim 2. Hence, it is immediate to check that $\mathcal{I}$ cannot be a copy of {}×$\{\emptyset\}\times\mathrm{Fin}$. This concludes the proof. ∎
Remark 3.4.
Pick an infinite set $T\subseteq\omega$ which is not cofinite and define $\mathcal{J}:=\{S\subseteq\omega:S\cap T\in\mathrm{Fin}\}$ (hence, $\mathcal{J}$ is an isomorphic copy on $\omega$ of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$). Pick also an ideal $\mathcal{I}$ on $\omega$ such that $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$. Thanks to Theorem 2.4, there exists an infinite matrix $A\in(\ell_{\infty},\ell_{\infty})$ such that equality (2) holds. Now, as it has been observed in the proof of the If part of Theorem 2.5, we have also $\mathcal{J}\leq_{\mathrm{RK}}\mathrm{Fin}$. Hence, with the same argument, there exists a matrix $B\in(\ell_{\infty},\ell_{\infty})$ such that $\mathrm{core}_{B\bm{x}}(\mathrm{Fin})=\mathrm{core}_{\bm{x}}(\mathcal{J})$ for all sequences $\bm{x}\in\ell_{\infty}$. In addition, by the proof of Theorem 2.4, it is possible to assume that each row of $B$ contains a single $1$. Set $C:=BA$ (observe that each entry of $C$ is well defined) and note that, if $\bm{x}\in\ell_{\infty}$ then $C\bm{x}=B(A\bm{x})$ is bounded, hence $C\in(\ell_{\infty},\ell_{\infty})$. It follows also that
$\forall\bm{x}\in\ell_{\infty},\quad\mathrm{core}_{C\bm{x}}(\mathrm{Fin})=\mathrm{core}_{A\bm{x}}(\mathcal{J})=\mathrm{core}_{\bm{x}}(\mathcal{I}).$
This does not contradict the claim of Theorem 2.5. In fact, we claim that, if $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$ then either $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}$ is an isomorphic copy of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$.
For, let $h:\omega\to\omega$ be a map which witnesses $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$. Observe that $W:=h[T]\in\mathcal{I}^{+}$ since $T\in\mathcal{J}^{+}$, hence $W$ is an infinite set. Considering that $h^{-1}[F]\in\mathrm{Fin}$ if and only if $F\in\mathrm{Fin}$ for each $F\subseteq\omega$, we obtain that
$\begin{split}\forall S\in\omega,\quad S\in\mathcal{I}&\,\,\Longleftrightarrow\,\,h^{-1}[S]\in\mathcal{J}\,\,\Longleftrightarrow\,\,h^{-1}[S]\cap T\in\mathrm{Fin}\\ &\,\,\Longleftrightarrow\,\,h^{-1}[S\cap W]\in\mathrm{Fin}\,\,\Longleftrightarrow\,\,S\cap W\in\mathrm{Fin}.\end{split}$
To sum up, $W$ is an infinite set and =$\mathcal{I}=\{S\subseteq\omega:S\cap W\in\mathrm{Fin}\}$. Therefore either $\mathcal{I}=\mathrm{Fin}$ (in the case where $W$ is cofinite) or $\mathcal{I}$ is isomorphic to $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ (in the opposite case).
4. Acknowledgments
The author is grateful to an anonymous referee for a careful reading of the manuscript and several useful observations.